### v1

@@ 4.4 Ablation Study
Two ablation studies were conducted to validate the effectiveness of the CA module and RICA block, as well as the effectiveness of the weighted Focal LogDice loss in segmenting small and sparse lesions with high class imbalance.
Figure 4: Visualization of the attention maps generated by U-Net, CAU-Net, and RICAU-Net using Grad-CAM [30] with respect to the ground truth and prediction overlaid on the input image. The heatmap visualizes the intensity of the model’s attention to specific classes. The calcium lesion with a red heatmap area showed stronger attention in the model’s classification. The input image with similar per-lesion Dice scores for all three models was used to generate the attention maps. The per-lesion Dice scores of each model are as follows: LM: 100%, LAD: 99.82%, LCX: 100% for U-Net; LM: 100%, LAD: 99.46%, LCX: 100% for CAU-Net; LM: 100%, LAD: 99.64%, LCX: 100% for RICAU-Net. The per-lesion attention scores (AS) of each model are as follows: LM: 75%, LAD: 85.14%, LCX: 90% for U-Net; LM: 80.77%, LAD: 89.42%, LCX: 100% for CAU-Net; LM: 82.69%, LAD: 88.53%, LCX: 100% for RICAU-Net.
4.4.1 Effectiveness of the CA Module and RICA Block.
To validate the effectiveness of the CA module and RICA block, three architectures were compared: the vanilla U-Net, Coordinate Attention U-Net (CAU-Net), and RICAU-Net. CAU-Net has the same encoder as the vanilla U-Net and the same decoder as RICAU-Net (Fig. 1). All three architectures were trained using the same implementation details as RICAU-Net. Per-lesion Dice scores of the three models are shown in Tab. 2. For CAC in LM, CAU-Net exhibited a 3.37% higher Dice score compared to U-Net. RICAU-Net exhibited higher Dice scores for all four lesions compared to CAU-Net and UNet. RICAU-Net achieved a Dice score of 59.35 for CAC in LM, which was 2.96% and 6.33% higher than that of CAU-Net and U-Net, respectively.
In addition, Gradient-weighted Class Activation Mapping (Grad-CAM) [30] was used to visualize the attention maps generated by the three models. Grad-CAM is a technique that quantifies the contribution level of each feature map for a specific class to the classification decision of a model. The contribution levels are quantified in the range of [0, 1] and displayed on the image using a heatmap. Red areas or larger values indicate a higher contribution to the model’s decision. This indicates that the model pays more attention to that area when classifying them into a specific class. On the other hand, blue areas indicate that the contribution of the area to the decision is less important. As shown in Fig. 4, RICAU-Net demonstrated higher attention levels when making decisions for CAC in LM, LAD, and LCX compared to U-Net and CAU-Net, even though all three models showed similar per-lesion Dice scores. To quantify the contribution level in relation to the prediction, the per-lesion attention score (AS) was calculated as the number of pixels with a contribution level greater than 0.6 divided by the total number of predicted pixels. The per-lesion attention scores are presented in Fig. 4. RICAU-Net achieved the highest LM AS of 82.69%, which was 1.92% and 7.69% higher than that of CAU-Net and U-Net, respectively. For CAC in LCX, RICAU-Net and CAU-Net achieved the same per-lesion Dice score of 100% and AS of 100%. However, the attention map generated by RICAU-Net displayed larger areas filled with a darker red color compared to CAU-Net. When comparing the attention maps and attention scores produced by CAU-Net and U-Net, it is significant that CAU-Net outperformed U-Net by achieving higher attention scores, as evidenced by its darker red-colored attention maps. Therefore, this ablation study demonstrated that the implementation of the CA module and the RICA block was effective in segmenting small and sparse lesions. The impact was particularly notable for CAC in LM, which faced a high class-imbalance problem.
Table 2: Per-lesion Dice scores for U-Net, CAU-Net, and RICAU-Net. All the architectures were trained under identical conditions using the weighted Focal LogDice loss function. The highest Dice score in each lesion is in bold.
Method Per-lesion Dice Scores (%)
LM LAD LCX RCA
U-Net 53.02 93.62 84.13 91.37
CAU-Net 56.39 92.94 85.02 91.96
RICAU-Net (ours) 59.35 94.11 87.03 92.89
Table 3: Per-lesion Dice scores of RICAU-Nets with four different loss functions. Four RICAU-Nets were trained using different loss functions. Other conditions were identical. The highest Dice score in each lesion is in bold.
@@
Weighted Focal 56.63 92.85 84.57 90.42
Weighted Focal Dice 55.48 93.68 85.00 91.42
Weighted Focal LogDice (ours) 59.35 94.11 87.03 92.89
4.4.2 Effectiveness of the Weighted Focal LogDice Loss.
To validate the effectiveness of the weighted Focal LogDice loss in segmenting small and sparse lesions, RICAU-Net was compared with four different loss functions: Cross-entropy loss, weighted Focal loss, weighted Focal Dice loss, and weighted Focal LogDice loss. All the loss functions were trained using the identical implementation details as RICAU-Net. Per-lesion Dice scores of the four different loss functions with RICAU-Net are shown in Tab. 3. RICAU-Net with the proposed loss function achieved the highest Dice scores for all four lesions, and a significant improvement in the Dice score for CAC in LM was observed. The proposed loss function showed a 10.51% higher Dice score for CAC in LM compared to the lowest score, while the Dice scores for CAC in LAD, LCX, and RCA increased by 1.48%, 2.46%, and 2.47% compared to the lowest scores, respectively. This demonstrated that the weighted Focal LogDice loss function is useful in the segmentation of small and sparse lesions, especially for classes experiencing a high class-imbalance problem.
@@ 5 Conclusion
In this study, we proposed a deep learning architecture called RICAU-Net and a customized combo loss function named the weighted Focal LogDice loss for the lesion-specific CAC segmentation task. This task generally suffers from a high class-imbalance problem with small and sparse lesion datasets. To overcome the problems, the coordinate attention was utilized in the form of RICA blocks in the encoder and as CA modules in the decoder of the proposed architecture. The two ablation studies emphasized the importance of the CA module, the RICA block, and the proposed loss function in segmenting small and sparse lesions, particularly for CAC in LM which was small and the scarcest in our dataset. The Grad-CAM attention maps produced by RICAU-Net exhibited a darker red color with higher attention scores on the calcium lesions compared to those of CAU-Net. This trend persisted when comparing CAU-Net to U-Net. This indicates that the presence of RICA blocks and CA modules made the architecture pay more attention to the calcium lesions. When the proposed method was compared with other methods, our method exhibited the highest per-lesion Dice scores for all four lesions. Additionally, it also showed the highest per-lesion precision for CAC in LM and LCX, and the highest per-lesion recall for CAC in LAD and RCA.
This study still has two limitations. First, the complex cases were not adequate in all datasets. Considering the complex cases with calcium lesions in more than three coronary arteries, there were only 41, 6, and 9 complex images in the training, validation, and test sets, respectively. And not all the complex case images showed the lesions coexisting as a cluster. Additionally, there was no image showing CAC in LM, LAD, LCX, and RCA simultaneously. Second, the ground truth only had six classes: background, bone, CAC in LM, LAD, LCX, and RCA. The calcium lesions in the aorta, ramus intermedius artery, aortic valve, mitral valve, and stent were all annotated and included as bone. Hence, if more datasets with distinct annotations for the mentioned classes were provided, a higher performance of the proposed method could be anticipated.
In conclusion, we propose an end-to-end deep learning algorithm and a customized combo loss function for lesion-specific CAC segmentation using non-contrast cardiac CT scans. This algorithm can be further utilized to streamline the manual process of identifying lesion-specific calcium plaques and calculating lesion-specific calcium scores by cardiologists.
5.0.1 Acknowledgements.
This work was supported by National Medical Research Council (NMRC) of Singapore (NMRC/CG2/001a/2021-NHCS, CG21APR1006, TA21nov-0001). The computational work for this article was fully performed on resources of the National Supercomputing Centre (NSCC), Singapore (https://www.nscc.sg).

### v2

@@ 3.6 Ablation Study
An ablation study was conducted to validate the effectiveness of the weighted Focal LogDice loss in segmenting small and sparse lesions with high class imbalance. Four different loss functions were compared using RICAU-Net: Cross-entropy loss, weighted Focal loss, weighted Focal Dice loss, and weighted Focal LogDice loss. All the loss functions were trained using the identical implementation details as RICAU-Net. Per-lesion Dice scores of the four different loss functions with RICAU-Net are shown in Table 2. RICAU-Net with the proposed loss function achieved the highest Dice scores for all four lesions, and a significant improvement in the Dice score for CAC in LM was observed. The proposed loss function showed a 10.51% higher Dice score for CAC in LM compared to the lowest score. This demonstrated that the weighted Focal LogDice loss function is useful in the segmentation of small and sparse lesions, especially for classes experiencing a high class-imbalance problem.
Table 2: Per-lesion Dice scores of RICAU-Nets with four different loss functions. All the loss functions were weighted except for the cross-entropy loss. The highest per-lesion Dice scores are in bold.
@@
Focal 56.63 92.85 84.57 90.42
Focal Dice 55.48 93.68 85.00 91.42
Focal LogDice (ours) 59.35 94.11 87.03 92.89
@@ 4 Conclusion
In this study, we proposed a deep learning architecture called RICAU-Net and a customized combo loss function named the weighted Focal LogDice loss for the lesion-specific CAC segmentation using non-contrast cardiac CT scans that generally suffers from a high class-imbalance problem with small and sparse lesion datasets. When the proposed method was compared with other methods used in medical image segmentation tasks, our method exhibited the highest per-lesion Dice scores for all four lesions. The ablation study emphasized the importance the proposed loss function in segmenting small and sparse lesions, particularly for CAC in LM which was small and the scarcest in our dataset. This algorithm can be further utilized to streamline the manual process of identifying lesion-specific calcium plaques and calculating lesion-specific calcium scores by cardiologists.
5 Compliance with ethical standards
This research study was conducted retrospectively using human subject data. Ethical approval was received from the Institutional Review Board of SingHealth (IRB 2022-2521).
6 Acknowledgments
This work was supported by National Medical Research Council (NMRC) of Singapore (grant numbers: NMRC/CG2/001a/2021-NHCS, CG21APR1006, TA21nov-000), SingHealth-Duke-NUS ACP Philanthropy grant (HRDUK230600) and SingHealth-Duke-NUS ACP NCRS grant (07/FY2021/P2/12-A93). L.Z. is supported by NMRC under its Clinician-Innovator Award Senior Investigator (CIASI24jan-0001) and Health Service Research Award (MOH-000358), SingHealth Duke-NUS Academic Medical Centre AM strategic fund award (07 FY2023 HTP P2 15-A2) and Industry Alignment Fund Pre-Positioning (IAFPP award) (H20c6a0035). The computational work for this article was fully performed on resources of the National Supercomputing Centre (NSCC), Singapore (https://www.nscc.sg).